# SiLVR: Scalable Lidar-Visual Radiance Field Reconstruction with Uncertainty Quantification

### v1

@@ III-B Bayesian Interpretation of the NeRF training
Here, the posterior $p(\theta|\mathcal{D})$ is used for the uncertainty estimation described later in Sec. IV-C
@@ IV-A Geometric Constraints from Lidar Measurements
Image-based 3D reconstruction with NeRF becomes challenging when a surface has uniform texture and limited multi-view constraints. Lidar measurements are complementary as they can provide accurate depth measurements in such scenarios. In our work, we incorporate the lidar measurements directly into the NeRF optimization. Specifically, we project the lidar point cloud from VILENS-SLAM’s pose graph onto the image plane using the camera intrinsics and lidar-camera extrinsics (described in Sec. V-B) to form a depth image. We denote the lidar depth images as $\mathcal{D}_{d}$. Each frame of the lidar point cloud is motion-undistorted to the nearest image’s timestamp, and hence synchronised with the image data. Example overlays can be found in Fig. 4.
@@ IV-A1 Lidar Depth Constraints
We follow the depth regularisation approach proposed by DS-NERF [10] for RGB-D cameras, which minimises the Kullback-Leibler (KL) divergence between a (narrow) normal distribution around the lidar depth-measurement $\mathbf{D}$ and the rendered ray distribution $h(t)$ from the NeRF model:
$\mathcal{L}_{\text{Depth-KL}}=\sum_{\mathbf{r}\in\mathcal{D}_{d}}\mathrm{KL}[\mathcal{N}(\mathbf{D},\hat{\sigma})\|h(t)]$ (7)
During training, we adopt a coarse-to-fine approach by gradually reducing the covariance of the target Normal distribution. Then, the effect of this loss function is that the surface is encouraged to be close to a delta function.
@@ IV-A2 Surface Normal Constraints from Lidar Measurements
While the depth loss improves 3D reconstruction, we found that the surface it produces will still contain wavy artefacts in regions where it is expected to be smooth (see Fig. 3). To mitigate this, we regularise the surface normal of the NeRF with lidar data. For each point $\mathbf{p}$ in the radiance field $R$, its surface normal can be computed as the negative gradient of the volume density $\sigma$. To obtain the training labels for the surface normal, we use the lidar range image and estimate the surface normals by local plane fitting. The surface normals are projected onto the image plane to generate the surface normal images, denoted as $\mathcal{D}_{n}$ similar to the lidar depth images. Then, we introduce an additional surface normal regularisation loss in the NeRF training, inspired by [11]:
@@
Compared to learning-based surface normal estimation from the camera image used in [11], our surface normal is estimated from the 3D lidar point cloud and does not suffer generalisation issues. The effect of the surface normal regularisation can be seen in Fig. 3.
@@ IV-C Epistemic Uncertainty of the NeRF reconstruction
We aim to obtain a explicit metric of uncertainty of our NeRF reconstruction. Particularly, following Sec. III-B, we aim to quantify the epistemic uncertainty that is directly related to the covariance of the posterior distribution $p(\theta|\mathcal{D})$ using the Laplace approximation. We first describe the reformulation of the parameters $\theta$, and then we derive the epistemic uncertainty estimate using the approximation.
@@ IV-C1 Perturbation Field Reformulation
While a NeRF representation presents convenient advantages for scene compression, its parameters $\theta$ do not directly correspond to the 3D scenes. A change of one parameter in the MLP could change the whole radiance field, and it is difficult to obtain uncertainty for a specific 3D location. This is in contrast to other approaches such as 3D Gaussian Splatting [4], where the parameters have a direct representation in the world. This introduces additional challenges when estimating the uncertainty of the parameters.
To obtain the spatial uncertainty of a NeRF, we adopt the perturbation field formulation introduced in BayesRays [15]. Specifically, we construct the perturbation field $\mathcal{P}$ of every 3D coordinate $\mathbf{x}$. We define the perturbation field as $\mathcal{P}_{\theta_{N}}(\mathbf{x}):\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$, where $\theta_{P}$ denotes the parameters in the perturbation field. A 3D coordinate’s perturbation can be obtained using trilinear interpolation:
$\mathcal{P}_{\theta_{N}}(\mathbf{x})=\text{Trilinear}(\mathbf{x},\theta_{N})$ (10)
@@ IV-C2 Epistemic Uncertainty Estimation using Laplace Approximation
We estimate the epistemic uncertainty of the NeRF reconstruction by estimating the covariance of the posterior $p(\theta|\mathcal{D})$. Using Laplace approximation, we estimate the otherwise intractable posterior distribution as a Gaussian function, and then use its covariance as the estimated uncertainty of our reconstruction. Our derivation follows the presentation by Daxberger et al. [14].
$\mathcal{L}(\mathcal{D};\theta)\approx\mathcal{L}(\mathcal{D};\hat{\theta})+\frac{1}{2}(\theta-\hat{\theta})^{\top}\mathbf{H}(\theta-\hat{\theta}),$ (11)
where $\mathbf{H}=\nabla_{\theta}^{2}\mathcal{L}(\mathcal{D};\theta)|_{\hat{\theta}}$ is the Hessian matrix at the MAP estimate $\hat{\theta}$. Here, the first-order term $\nabla_{\theta}\mathcal{L}(\mathcal{D};\theta)|_{\hat{\theta}}^{\top}(\theta-\hat{\theta})$ does not appear as it is zero at the MAP estimate.
Substituting the approximation in Eq. (11) into Eq. (6), we obtain:
[…]
where $D$ denotes the dimensionality of the parameters $\theta$.
Then, we substitute the Taylor expansion Eq. (11) and expression of the normalization constant Eq. (12) into the posterior Eq. (5):
$p(\theta|\mathcal{D})\approx\frac{(\det\mathbf{H})^{\frac{1}{2}}}{(2\pi)^{\frac{D}{2}}}\exp\left(-\frac{1}{2}(\theta-\hat{\theta})^{\top}\mathbf{H}(\theta-\hat{\theta})\right)$ (13)
which turns out to be a Gaussian distribution $\mathcal{N}(\theta;\hat{\theta},\mathbf{\Sigma})$ with mean $\hat{\theta}$ and covariance $\mathbf{\Sigma}=\mathbf{H}^{-1}$.
By now, we have shown that the posterior can be approximated as a Gaussian distribution. The mean of the Gaussian $\hat{\theta}$ is the MAP estimate of the parameters $\theta$. Here, the parameters $\theta$ that we are estimating are the perturbation field $\theta_{P}$ introduced in Sec. IV-C1. Assuming that the NeRF reconstruction has converged to local minima after training, small perturbation should not make the reconstruction, and hence the MAP estimate of the perturbation field is $\mathbf{0}$ (no perturbation). Then, the major computation needed is to determine the covariance, or the inverse of the Hessian. If we assume the prior is a zero-mean Gaussian $p(\theta)=\mathcal{N}(\theta;\mathbf{0},\gamma^{2}\mathbf{I})$, then the full Hessian at the location of the MAP estimate is
@@
Since the Fisher information matrix is still expensive to compute, we can make a further assumption that each parameter (perturbation field) is independent of each other, and use a diagonal approximation to the Hessian:
@@
The Hessian matrix is initialised as all zeroes, i.e. infinite covariance. As a result, the parameters that are not involved in the outputs (the rendered pixels or depth) will have very high epistemic uncertainty because changing them will not change the outputs and the training loss. Since the outputs are rendered according to the rays from the training images, the unobserved regions can be identified as having very high uncertainty and can be filtered out. Identifying unobserved areas (similar to the unknown space in occupancy mapping) is particularly important for NeRF, as the underlying MLP can output arbitrary colour and density in some locations, leading to “hallucinated” reconstruction points.
IV-C3 Visual and Depth Uncertainty
The NeRF model is trained with a total training loss function $\mathcal{L}(\mathcal{D;\hat{\theta}})$ that contains the photometric loss $\mathcal{L}_{Colour}$ and the depth loss $\mathcal{L}_{Depth}$. This means that the Jacobian $\mathbf{J}$ can be decoupled into a colour component $\mathbf{J}_{Colour}$ and depth component $\mathbf{J}_{Depth}$, from which we can then approximate the Hessian that corresponds only to the visual information $\mathbf{H}_{Colour}$, and the Hessian that corresponds to only to the lidar depth information $\mathbf{H}_{Depth}$. Therefore, we can compute the epistemic uncertainty for each observation modality by changing the loss function. Note that the total training loss function contains other terms including the surface normal loss $\mathcal{L}_{Normal}$. In our study, we focus on just the photometric loss $\mathcal{L}_{Colour}$ and depth loss $\mathcal{L}_{Depth}$. This is because these two losses are the dominant components of the total loss $\mathcal{L}(\mathcal{D;\hat{\theta}})$ (with our weighting coefficients for each loss chosen).
The nature of the different sensor modalities leads to different uncertainty characteristics. The visual uncertainty captures areas that can be geometrically perturbed while not changing the colour. As later shown in Fig. 8, “low” visual uncertainty corresponds to distinct visual features and high-frequency areas. “Higher” visual uncertainty typically corresponds to areas with uniform texture, since perturbing a point in an area with similar colours can lead to small changes in the rendered colour. Here, the characteristics of the visual uncertainties are similar to those in SfM where visual features are used as landmarks for Bundle Adjustment whereas uniform texture areas are often not mapped.
In comparison, low lidar depth uncertainty often appears in regions with abundant lidar observation—whether there are visual features or not. This means a road surface with uniform texture can have lower lidar depth uncertainty but higher visual uncertainty. Low lidar depth uncertainty is also observed at object boundaries, since perturbing that point can lead to a drastic change in the depth (from foreground depth to background depth). Regions with high lidar depth uncertainty are often areas where there are fewer lidar observations, such as the sky, distant regions that are beyond the lidar’s sensing range, and regions outside the lidar’s Field-of-View.
@@ IV-D1 Refining Lidar SLAM trajectory with Bundle Adjustment
Providing the NeRF reconstruction method with accurate camera poses is crucial as their accuracy directly impacts the fidelity of the reconstructed model. A popular approach used in most NeRF works is to estimate camera poses using (offline) Structure-from-Motion methods such as COLMAP [1]. However, we observed the following limitations when testing COLMAP: (1) long computation times, especially for large image collections collected spanning a long trajectory (e.g., 3000 images can take more than 1 hour (as shown in Tab. IV), and (2) inability to register all frames into one global map when there is limited visual overlap between the images. These issues undermine the goal of building complete, large-scale, globally consistent maps.
@@
[…]
@@ IV-D2 Submapping of Pose Trajectory
We formulate the clustering problem as a graph partitioning problem, where each node is an image, and the edges between the nodes are weighted by a similarity score. We measure the similarity between two images using a co-visibility metric. Two images are co-visible if some feature points on one image can be viewed from the other image. Specifically, the co-visibility metric for an image pair is computed as the number of visual feature points computed by COLMAP that appear in both images. After constructing the graph, we use the Normalised Cuts algorithm [18] to obtain a partitioning which minimally breaks edges, i.e. to remove the link between unrelated nodes. In practice, this means that images that are co-visible are grouped together, and image pairs that have less visual overlap are identified as the submap boundary.
Once we divide the full map into submaps as a set of clustered images, we independently train each NeRF submap. After training, we compute the epistemic uncertainty of the reconstruction. We can export a point cloud by rendering colour and depth using the training data rays, and we filter points that have high uncertainty.
Figure 4: Sample Data from our diverse robotic datasets. Here each image is overlaid with projected lidar point cloud to demonstrate the accuracy of the sensor calibration.
@@ V-A Hardware and Datasets
We evaluate our methods using a custom perception unit called Frontier shown in Fig. 2. It includes three 1.6 MP fisheye colour [^2] To produce RGB images, we debayer and white-balance the raw bayered images using https://github.com/leggedrobotics/raw_image_pipeline Alphasense cameras (from Sevensense Robotics AG) on 3 sides of the device, as well as a synchronised IMU. The 3-camera setup enables omnidirectional NeRF mapping from a single walking pass through a test site. We installed a Hesai QT64 lidar (104° Field-of-View, 60 metres maximum range) on top of the device. We deployed the Frontier in different modes—onboard a legged robot (Boston Dynamics Spot), a drone (DJI M600, with the system described in [59]), or simply handheld (using the Oxford Spires dataset [19]. Some data is collected with the Frontier mounted on a human operator’s backpack).
@@
[…]

### v3

@@ III-C Laplace Approximation
[…]
Substituting the approximation in Eq. (7) into Eq. (6), we obtain:
$Z$ $=\int p(\mathcal{D}\mid\theta)p(\theta)\,d\theta$ (8)
$=\int\exp(-\mathcal{L}(\mathcal{D};\theta))d\theta$
$\approx\exp(-\mathcal{L}(\mathcal{D};\hat{\theta}))\int\exp\left(-\frac{1}{2}(\theta-\hat{\theta})^{\top}\mathbf{H}(\theta-\hat{\theta})\right)d\theta$
where $D$ denotes the dimensionality of the parameters $\theta$.
We then substitute the Taylor expansion Eq. (7) and expression of the normalization constant Eq. (8) into the posterior Eq. (5):
[…]
which corresponds to a Gaussian distribution $\mathcal{N}(\theta;\hat{\theta},\mathbf{\Sigma})$ with mean $\hat{\theta}$ and covariance $\mathbf{\Sigma}=\mathbf{H}^{-1}$.
By using Laplace approximation, the uncertainty estimation problem can be formulated as estimating the Hessian matrix $\mathbf{H}$ at the MAP estimate $\hat{\theta}$. We describe how we apply this to the NeRF reconstruction problem in Sec. IV-C2, and techniques to further approximate the Hessian matrix in Sec. IV-C3.
@@ IV Method
In this section, we present SiLVR, a 3D reconstruction system based on a NeRF representation. An overview of our system is presented in Fig. 2. SiLVR takes in as input synchronised triplets of a camera image, a lidar depth image and the corresponding surface normals, as well as the sensor trajectory estimated by an online Lidar SLAM system. This trajectory is refined and partitioned into submaps. For each submap, we train a NeRF which combines both vision and lidar data. After training the NeRF, SiLVR estimates the epistemic uncertainty of the NeRF. The final point cloud reconstruction is extracted from the NeRF based on rendered depth, and filtered using the per-point uncertainty estimates.
We describe our approach to extend the NeRF formulation to include lidar data in Sec. IV-A, and how to remove the sky reconstruction in Sec. IV-B. We then introduce the epistemic uncertainty estimation method in Sec. IV-C. Finally, we present how SiLVR achieves scalability by partitioning large-scale poses into submaps in Sec. IV-D.
@@ IV-A Geometric Constraints from Lidar Measurements
Image-based 3D reconstruction with NeRF becomes challenging when a surface has a uniform texture and limited multi-view constraints. Lidar measurements are complementary as they can provide accurate depth measurements in such scenarios. In our work, we incorporate the lidar measurements directly into the NeRF optimization. Specifically, we project the lidar point cloud from VILENS-SLAM’s pose graph onto the image plane using the camera intrinsics and lidar-camera extrinsics (described in Sec. V-B) to form a depth image. We denote the lidar depth images as $\mathcal{D}_{d}$. Each frame of the lidar point cloud is motion-undistorted  [^1] See implementations at https://github.com/ethz-asl/lidar_undistortion to the nearest image’s timestamp, and hence synchronised with the image data. Example overlays can be found in Fig. 4.
@@ IV-A1 Lidar Depth Constraints
We follow the depth regularisation approach proposed by DS-NeRF [16] for RGB-D cameras. We define the rendering weight distribution $w(t)$ as a discrete probability distribution, with probabilities given by $w_{i}$, the weights of the ray samples defined in Eq. (2). The depth regularisation term is obtained by minimising the Kullback-Leibler (KL) divergence between a narrow normal distribution centred at the lidar depth measurement $\mathbf{D}$ and the rendering weight distribution $w(t)$:
$\mathcal{L}_{\text{Depth-KL}}=\sum_{\mathbf{r}\in\mathcal{D}_{d}}\mathrm{KL}[\mathcal{N}(\mathbf{D},\hat{\sigma})\|w(t)]$ (10)
During training, we apply this regularisation using a coarse-to-fine schedule by progressively reducing the covariance $\hat{\sigma}$ of the target Normal distribution. This encourages the rendering weight distribution $w(t)$ to approach a Dirac delta function. As a result, the density along the ray is encouraged to be concentrated near the lidar depth.
@@ IV-A2 Surface Normal Constraints from Lidar Measurements
While the depth loss improves 3D reconstruction, we found that the surface it produces will still contain wavy artefacts in regions where it is expected to be smooth (see Fig. 3). To mitigate this, we regularise the surface normal of the NeRF with lidar data. For each point $\mathbf{p}$ in the radiance field $R$, its surface normal can be computed as the negative gradient of the volume density $\sigma$. To obtain the training labels for the surface normal, we use the lidar range image and estimate the surface normals by local plane fitting. The surface normals are projected onto the image plane to generate the surface normal images, denoted as $\mathcal{D}_{n}$, similar to the lidar depth images. Then, we introduce an additional surface normal regularisation loss in the NeRF training, inspired by [17]:
@@
Compared to learning-based surface normal estimation from the camera image used in [17], our surface normal is estimated from the 3D lidar point cloud and does not suffer from generalisation issues. The effect of the surface normal regularisation can be seen in Fig. 3.
@@ IV-B Sky Segmentation
Combining the loss terms introduced in Eq. (3), Eq. (10), Eq. (11) Eq. (12), our overall training loss is
$\mathcal{L}=\mathcal{L}_{\text{Colour}}+\mathcal{L}_{\text{Depth-KL}}+\mathcal{L}_{\text{Normal }}+\mathcal{L}_{\text{Sky}}$ (13)
@@ IV-C Epistemic Uncertainty of the NeRF reconstruction
We aim to obtain an explicit metric of uncertainty of our NeRF reconstruction. Particularly, following Sec. III-B, we aim to quantify the epistemic uncertainty that is directly related to the covariance of the posterior distribution $p(\theta|\mathcal{D})$ using the Laplace approximation. We first describe the reformulation of the parameters $\theta$, and then we derive the epistemic uncertainty estimate using the approximation.
@@ IV-C1 Perturbation Field Reformulation
While a NeRF representation presents convenient advantages for scene compression, its parameters $\theta$ do not directly correspond to the 3D scenes. A change of one parameter in the MLP could change the whole radiance field, and it is difficult to obtain uncertainty for a specific 3D location. This is in contrast to other approaches, such as 3D Gaussian Splatting [4], where the parameters have a direct representation in the world. This introduces additional challenges when estimating the uncertainty of the parameters.
To obtain the spatial uncertainty of a NeRF, we adopt the perturbation field formulation introduced in BayesRays [22]. Specifically, we construct the perturbation field $\mathcal{P}$ of every 3D coordinate $\mathbf{x}$. We define the perturbation field as $\mathcal{P}_{\theta_{P}}(\mathbf{x}):\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$, where $\theta_{P}$ denotes the parameters in the perturbation field. A 3D coordinate’s perturbation can be obtained using trilinear interpolation:
$\mathcal{P}_{\theta_{P}}(\mathbf{x})=\text{Trilinear}(\mathbf{x},\theta_{P})$ (14)
When estimating the uncertainty, we modify the volume rendering procedure by adding the perturbation field to the point samples $\mathbf{p}_{i}$ along the ray $\mathbf{r}$ to produce the new point sample $\mathbf{p}_{i}^{\prime}=\mathbf{p}_{i}+\mathcal{P}_{\theta_{P}}(\mathbf{p}_{i})$. The new point colour and volume density $(\mathbf{c}_{i}^{\prime},\sigma_{i}^{\prime})=R(\mathbf{p}^{\prime},\mathbf{d})$, and the new pixel colour $\hat{\mathbf{C}^{\prime}}(\mathbf{r})$ can then be computed using Eq. (1) and Eq. (2).
@@ IV-C2 Epistemic Uncertainty Estimation using Laplace Approximation
We estimate the epistemic uncertainty of the NeRF reconstruction by estimating the covariance of the posterior $p(\theta|\mathcal{D})$. Using the Laplace approximation technique described in Sec. III-C, we estimate the otherwise intractable posterior distribution as a Gaussian function, and then use its covariance as the estimated uncertainty of our reconstruction.
Specifically, the mean of the Gaussian $\hat{\theta}$ is the MAP estimate of the parameters $\theta$. Here, the parameters $\theta$ that we are estimating are the perturbation field $\theta_{P}$ introduced in Sec. IV-C1. Assuming that the NeRF reconstruction has converged to a local minima after training, a small perturbation should not cause the reconstruction to deteriorate, and hence the MAP estimate of the perturbation field is $\mathbf{0}$ (no perturbation). Then, the major computation needed is to determine the covariance, or the inverse of the Hessian. If we assume the prior is a zero-mean Gaussian $p(\theta)=\mathcal{N}(\theta;\mathbf{0},\gamma^{2}\mathbf{I})$, then the full Hessian at the location of the MAP estimate is
@@ IV-C3 Approximation of the Hessian Matrix
Since the Fisher information matrix is still expensive to compute, we can make a further assumption that each parameter (perturbation field) is independent of the others, and use a diagonal approximation to the Hessian:
@@
The Hessian matrix is initialised as all zeroes; i.e., infinite covariance. After accumulating the gradients, the prior term $\gamma^{-2}$ in Eq. (18) ensures that the final Hessian matrix is positive definite and the covariances are bounded. As a result, the parameters that are not involved in the outputs (the rendered pixels or depth) will have very high epistemic uncertainty because changing them will not change the outputs and the training loss. Since the outputs are rendered according to the rays from the training images, the unobserved regions can be identified as having very high uncertainty and can be filtered out. Identifying unobserved areas (similar to the unknown space in occupancy mapping) is particularly important for NeRF, as the underlying MLP can output arbitrary colour and density in some locations, leading to “hallucinated” reconstruction points.
IV-C4 Uncertainty for Heterogeneous Sensors
[…]
The nature of the different sensor modalities leads to different uncertainty characteristics. The visual uncertainty captures areas that can be geometrically perturbed while not changing the colour. As later shown in Fig. 8, “low” visual uncertainty corresponds to distinct visual features and high-frequency areas. “High” visual uncertainty typically corresponds to areas with uniform texture, since perturbing a point in an area with similar colours can lead to small changes in the rendered colour. Here, the characteristics of the visual uncertainties are similar to those in SfM where visual features are used as landmarks for Bundle Adjustment, whereas uniform texture areas are often not mapped.
In comparison, low lidar depth uncertainty often appears in regions with abundant lidar observation—whether there are visual features or not. This means a road surface with uniform texture can have lower lidar depth uncertainty but higher visual uncertainty. Low lidar depth uncertainty is also observed at object boundaries, since perturbing that point can lead to a drastic change in the depth (from foreground depth to background depth). Regions with high lidar depth uncertainty are often areas where there are fewer lidar observations, such as the sky, distant regions that are beyond the lidar’s sensing range, and regions outside the lidar’s FOV.
@@ IV-D1 Refining Lidar SLAM trajectory with Bundle Adjustment
Providing the NeRF reconstruction method with accurate camera poses is crucial, as their accuracy directly impacts the fidelity of the reconstructed model. A popular approach used in most NeRF works is to estimate camera poses using (offline) Structure-from-Motion methods such as COLMAP [1]. However, we observed the following limitations when testing COLMAP: (1) long computation times, especially for large image collections collected spanning a long trajectory (e.g., 3000 images can take more than 1 hour (as shown in Tab. IV), and (2) inability to register all frames into one global map when there is limited visual overlap between the images. These issues undermine the goal of building complete, large-scale, globally consistent maps.
@@
After COLMAP computation, we rescale the trajectory using $\mathit{Sim}(3)$ Umeyama alignment to the Lidar-SLAM trajectory, so that the final trajectory is metrically scaled. This step is crucial because the lidar measurements used in Sec. IV-A1 are also metric. The depth regularisation cannot be used if the scale of the scene and the scale of the depth are not consistent.
@@ IV-D2 Submapping of Pose Trajectory
We formulate the clustering problem as a graph partitioning problem, where each node is an image, and the edges between the nodes are weighted by a similarity score. We measure the similarity between two images using a co-visibility metric. Two images are co-visible if some feature points on one image can be viewed from the other image. Specifically, the co-visibility metric for an image pair is computed as the number of visual feature points computed by COLMAP that appear in both images. After constructing the graph, we use the Normalised Cuts algorithm [25] to obtain a partitioning which minimally breaks edges; i.e., to remove the link between unrelated nodes. In practice, this means that images that are co-visible are grouped together, and image pairs that have less visual overlap are identified as the submap boundary.
Once we divide the full map into submaps as a set of clustered images, we independently train each NeRF submap. After training, we compute the epistemic uncertainty of the reconstruction. We can export a point cloud by rendering colour and depth using the training data rays, and we filter out points that have high uncertainty. Rendering at the submap intersections can be obtained by combining two submap renderings in the image space where the weights are determined based on the distances to the neighbouring submap, following Block-NeRF [13].
Figure 4: Sample Data from our diverse robotic datasets. Here, each image is overlaid with a projected lidar point cloud to demonstrate the accuracy of the sensor calibration.
@@ V-A Hardware and Datasets
We evaluate our methods using a custom perception unit called Frontier shown in Fig. 2. It includes three 1.6 MP fisheye colour [^3] To produce RGB images, we debayer and white-balance the raw bayered images using https://github.com/leggedrobotics/raw_image_pipeline Alphasense cameras (from Sevensense Robotics AG) on 3 sides of the device, as well as a synchronised IMU. The 3-camera setup enables omnidirectional NeRF mapping from a single walking pass through a test site. We installed a Hesai QT64 lidar (104° FOV, 60 metres maximum range) on top of the device. We deployed the Frontier in different modes—onboard a legged robot (Boston Dynamics Spot), a drone (DJI M600, with the system described in [65]), or simply handheld (using the Oxford Spires dataset [26]. Some data was collected with the Frontier mounted on a human operator’s backpack).
@@
We tested our method using data collected in the following sites: H B Allen Centre (HBAC), Fire Service College (FSC), Radcliffe Observatory Quarter (ROQ), and the Bodleian Library, all in Oxford, UK. The large-scale sites, ROQ and the Bodleian Library, cover areas of $\text{5,000}~\text{m}^{2}$ and $\text{15,000}~\text{m}^{2}$, respectively. The hardware details are in Tab. I, and some sample lidar-camera overlays are shown in Fig. 4.